prove that lim n->infinity [1/√2n-1 +1/√4n-2^2 +1/√6n-3^2 +.......+1/n]=π/2

1
Expert's answer
2022-01-11T00:00:11-0500

"\lim_{n\to \infty}\left[ \frac{1}{\sqrt{2n-1}}+\frac{1}{\sqrt{4n-2^2}}+\frac{1}{\sqrt{6n-3^2}}+\cdots +\frac{1}{n}\right]\\
\lim_{n\to \infty} \sum_{r=1}^n \frac{1}{\sqrt{2rn-r^2}}\\
\lim_{n\to \infty} \sum_{r=1}^n \frac{1}{n\sqrt{\frac{2r}{n}-(\frac{r}{n})^2}}\\"

Let "x=\frac{r}{n}, dx=\frac{1}{n}"

The limit becomes:

"\int_0^1\frac{1}{\sqrt{2x-x^2}}~dx\\
=\int_0^1\frac{1}{\sqrt{1-(x-1)^2}}~dx\\"

Let "x-1=t, dx=dt" . When "x=0, t=-1, x=1, t=0."

"\int_{-1}^0\frac{dt}{\sqrt{1-t^2}}=\sin^{-1}t \Biggr |_{-1}^0=\sin^{-1}0-\sin^{-1}(-1)=0-_-\frac{\pi}{2}=\frac{\pi}{2}"

Hence,

"\lim_{n\to \infty}\left[ \frac{1}{\sqrt{2n-1}}+\frac{1}{\sqrt{4n-2^2}}+\frac{1}{\sqrt{6n-3^2}}+\cdots +\frac{1}{n}\right]=\frac{\pi}{2}"